jigglypuff314 (jigglypuff314):

what is the derivative of: (sin(3x))^2 ?

OpenStudy (anonymous):

∫sin²(3x)dx Let u = 3x and du = 3dx, then... ∫sin²(3x)dx = (1/3)∫sin²udu sin²u = (1- cos2u)/2, trigonometric identity... (1/3)∫sin²udu = (1/3)∫[(1- cos2u)/2]du = (1/6)[∫(1)du - ∫(cos2u)du] =(1/6)[u - (1/2)sin2u] + c =(1/6)u - (1/12)sin2u + c As u = 3x, then... =(1/6)(3x) - (1/12)sin2(3x) + c =(1/2)x - (1/12)sin6x + c =D

jigglypuff314 (jigglypuff314):

... I still don't understand...

jigglypuff314 (jigglypuff314):

@piglet9 what am I doing wrong if I do deriv(sin(3x))^2 = 2(sin(3x))*(cos(3x))(3) = 6(sin(3x))(cos(3x))

OpenStudy (anonymous):

did I just do the integral... LOL

jigglypuff314 (jigglypuff314):

yeah... I don't think I've learned integrals...

OpenStudy (anonymous):

You're correct: 2sin(3x)∗cos(3x)∗3 =6∗sin(3x)cos(3x) =3(2sin(3x)cos(3x)) =3(sin(2∗3x)) =3sin(6x)

OpenStudy (anonymous):

Hope that helps :)

jigglypuff314 (jigglypuff314):

thanks :)